# Unified Optimization of Source Weights and Transfer Quantities in Multi-Source Transfer Learning: An Asymptotic Framework

### v1

@@ Abstract
Transfer learning plays a vital role in improving model performance in data-scarce scenarios. However, naive uniform transfer from multiple source tasks may result in negative transfer, highlighting the need to properly balance the contributions of heterogeneous sources. Moreover, existing transfer learning methods typically focus on optimizing either the source weights or the amount of transferred samples, while largely neglecting the joint consideration of the other. In this work, we propose a theoretical framework, Unified Optimization of Weights and Quantities (UOWQ), which formulates multi-source transfer learning as a parameter estimation problem grounded in an asymptotic analysis of a Kullback–Leibler divergence–based generalization error measure. The proposed framework jointly determines the optimal source weights and optimal transfer quantities for each source task. Firstly, we prove that using all available source samples is always optimal once the weights are properly adjusted, and we provide a theoretical explanation for this phenomenon. Moreover, to determine the optimal transfer weights, our analysis yields closed-form solutions in the single-source setting and develops a convex optimization–based numerical procedure for the multi-source case.
Building on the theoretical results, we further propose practical algorithms for both multi-source transfer learning and multi-task learning settings. Extensive experiments on real-world benchmarks, including DomainNet and Office-Home, demonstrate that UOWQ consistently outperforms strong baselines. The results validate both the theoretical predictions and the practical effectiveness of our framework.
@@ I Introduction
Transfer learning has emerged as a critical paradigm in modern machine learning, especially in scenarios where labeled data is scarce for the target task. By leveraging knowledge from related source tasks, transfer learning enables models to generalize more effectively and efficiently to the target task. In recent years, transfer learning has a wide range of applications. For example, in deep reinforcement learning, transfer learning plays a crucial role in accelerating policy learning and improving sample efficiency by transferring policies, representations, or experience [67]. In healthcare, transfer learning leverages large medical datasets to support rare disease diagnosis with limited clinical data [42]. In natural language processing, pretrained models built on large-scale transfer learning foundations have revolutionized performance across a wide range of downstream tasks [11]. While single-source transfer learning has been studied earlier, growing attention is now being paid to the multi-source transfer learning. In such scenarios, one key question is how to effectively use the data from the source tasks. For tasks that share more similarity with the target, its samples ought to be used more extensively, while for tasks that are less similar, its samples should be used with caution. In fact, transferred knowledge can even undermine the learning of target task, which is termed as negative transfer.[53] In multi-source transfer learning scenarios where source data are considered to be abundant with respect to target data, it is important to determine the effective amount of information to be transferred.
TABLE I: Comparison across matching-based transfer learning methods, based on whether they perform quantity optimization, weight optimization, are tailored to multi-sources, have task generality, have shot generality, and require target labels. The ‘✓’ represents obtaining the corresponding aspects, while ‘✗’ represents the opposite. Quantity optimization denotes the ability to determine how many source samples or data subsets should be transferred. Weight optimization denotes the ability to assign or learn weights for different source tasks or samples. Task generality denotes the ability to handle various target task types, and shot generality denotes the ability to avoid negative transfer in different target sample quantity settings, including few-shot and non-few-shot.
@@
From a statistical perspective, while leveraging additional data from source tasks can reduce estimation variance and improve generalization, heterogeneity between source and target tasks may introduce systematic bias, resulting in negative transfer. Moreover, in multi-source settings, different source tasks often exhibit substantially different levels of relevance to the target task, making effective transfer learning non-trivial. Existing approaches typically address this challenge from one of two separate perspectives. On the one hand, task- or source-weighting methods aim to control the influence of different sources by assigning importance weights, thereby mitigating bias caused by irrelevant or mismatched domains. In current research on task weighting, most existing methods focus on assigning weights to source models, while relatively few studies assign weights to samples for unified learning. Moreover, existing weighting frameworks typically assume that all source samples are utilized by default. On the other hand, sample selection methods address the problem of deriving the optimal transfer quantity and selecting the identity of source samples to be transferred. Considering source weighting and transfer quantity in isolation restricts the solution space and can lead to suboptimal solutions, whereas jointly optimizing them leads to more principled and effective solutions. The goal of this work is to establish a theoretical framework that determines both the optimal source weights and transfer quantities, which are directly assigned to individual samples for joint training. In Table I, we provide a detailed comparison of the proposed method with several existing methods.
@@
In this work, we develop a theoretical framework termed Unified Optimization of Weights and Quantities (UOWQ), to determine the optimal weights and quantities for source tasks in transfer learning via asymptotic analysis. Using the measure based on Kullback–Leibler (K-L) divergence, we formulate the multi-source transfer learning problem as a parameter estimation problem, which in turn derives an optimization problem of the source weights and transfer quantities. We address this problem through an asymptotic analysis, which leads to the conclusion that using all available source samples is always optimal once the weights are properly adjusted, and further provides a principled method to calculate the optimal source weights. Specifically, to determine the optimal transfer weight, our analysis provides closed-form solutions in the single-source case and develops a convex-optimization–based numerical procedure in the multi-source case. Furthermore, we investigate why using all samples remains optimal under the joint optimization of weights and quantities, and analyze how the optimal weights are influenced by domain discrepancy, model dimensionality, and sample size. Moreover, we introduce practical algorithms capable of supporting both multi-source transfer learning and multi-task transfer learning. Finally, we validate the effectiveness of our proposed method through experiments on real-world dataset.
@@
  • Theoretical framework. We introduce a mathematical framework, UOWQ, which formulates multi-source transfer learning as a parameter estimation problem and grounds it in an asymptotic analysis of a Kullback–Leibler divergence–based generalization error measure. The analysis yields closed-form solutions for the optimal transfer quantity and the optimal transfer weight in the single-source case, and proposes a convex-optimization-based solution method for the multi-source case.ă In particular, it provides a theoretical proof and interpretation for the fact that using all source samples is always optimal under the derived conditions.
  • Practical algorithm. We develop a practical weighting algorithm derived from the theory, which assigns sample-level source weights and can be directly applied to both multi-source transfer learning and multi-task learning. To cope with scarce target data and ensure robustness, the weights are updated dynamically during training. Specifically, the algorithms alternate between model optimization and source weight updates. At each stage, the model is trained using target data together with adaptively weighted source samples, while the source weights are updated based on the parameters of the current model state. This iterative procedure jointly refines the target model and the transfer strategy, and naturally extends to both learning scenarios.ă
@@
This work is an extension of our previous conference paper [58], and the additional contributions are summarized as follows. First, the journal version unifies the optimization of both the optimal transfer quantities and the optimal source weights, and provides the complete theoretical proofs. Second, we include new theoretical analyses, including comparisons with related theoretical frameworks and an explanation of why maximizing the total transfer quantity remains optimal under joint optimization. Third, on the algorithmic side, we extend the multi-source method to a full multi-task learning algorithm. Finally, the experimental section is substantially strengthened, covering larger datasets, more baselines, and extensive supplementary experiments, including ablation studies, robustness evaluations, analyses of computational efficiency, and compatibility with LoRA (Low-Rank Adaptation).
The remainder of this paper is organized as follows. Section II reviews the related work. Section III presents the problem formulation and introduces the asymptotic normality of the maximum likelihood estimator (MLE). Section IV develops the main theoretical results for computing the optimal source weights and introduces the practical algorithm. Section V empirically evaluates the proposed framework and validates the theoretical findings.
@@ II-A Multi-source Transfer Learning
Multi-source transfer learning (MSTL) leverages knowledge from multiple related source domains or tasks to improve performance on a target task, addressing challenges such as domain shift and negative transfer. Unlike single-source transfer, MSTL integrates diverse source knowledge through dynamic source weighting [28], shared feature space alignment [52], and ensemble- or meta-learning–based aggregation [66]. By emphasizing relevant sources while suppressing less informative ones, MSTL enhances generalization in heterogeneous settings, where effective learning requires balancing source contributions and mitigating distribution discrepancies.
@@ II-B Task weighting
Task weighting, also referred to as source weighting, is a representative class of matching-based approaches in multi-source transfer learning and multi-task learning (MTL), as it regulates the contribution of different tasks without explicitly aligning feature distributions. Its objective is to balance competing task objectives during joint optimization. Early approaches relied on static weighting, assigning fixed weights based on heuristics such as task priority or dataset size [2], but these methods lack adaptability to dynamic task interactions. Recent studies have therefore focused on dynamic task weighting strategies. For example, uncertainty-based weighting was introduced in [18], GradNorm balances tasks by normalizing gradient magnitudes [7], and DWA adjusts weights according to task learning speeds [30]. From a multi-objective optimization perspective, MGDA [41] formulates task weighting as a Pareto-optimal optimization problem.
Recent work has further extended task weighting to large-scale pretrained models, primarily through gradient-based matching mechanisms. For instance, [65] reweights tasks via gradient alignment, while [27] proposes CAGrad to mitigate conflicting gradients. Other studies explore meta-learning–based strategies [45] and task affinity to automate weighting. Despite these advances, task weighting remains challenged by nonstationary task relationships [33], scalability to large task sets, and the lack of strong theoretical guarantees [12, 4]. Benchmark datasets such as Meta-Dataset [48] have been introduced to facilitate standardized evaluation.
Most existing task-weighting methods operate at the model level, assigning scalar weights to source models in a mixture-of-experts fashion [34]. These approaches typically rely on task similarity [31], domain divergence [62], or validation performance [6]. In contrast, only a limited number of studies consider sample-level weighting, where source samples are reweighted and jointly trained with the target data in a unified framework [44, 61].
@@ II-C Transfer Learning Theory
Existing theoretical studies in transfer learning can generally be divided into two primary categories. The first category aims to define and quantify the similarity or relatedness between source and target tasks. Numerous metrics have been proposed in this context, including the $l_{2}$ distance [32], optimal transport cost [8], LEEP (Log Expected Empirical Prediction) [36], Wasserstein distance [44], OTCE (Optimal Transport-based Conditional Entropy) [5], LogME [56], NCE[46], GBC (Geometric-Based Correlation) [29], and maximal correlation-based measures [21]. For example, LEEP [36] provides a probabilistic framework for transferability estimation by comparing model predictions on the target task. Meanwhile, OTCE [5] leverages optimal transport theory to quantify transferability under distribution shifts. These measures provide principled ways to assess how well knowledge from a source task can potentially benefit a target task.
This work belongs to the second group, which is dedicated to developing theoretical measures that assess and bound the generalization error in transfer learning scenarios. Various generalization error measures have been proposed to guide the assignment of task or source weights, including mutual information  [1], ${\mathcal{X}}^{2}$-divergence [47], and $\mathcal{H}$-score [55]. Specifically, the $\mathcal{H}$-score facilitates the minimization of the bound of target generalization error by quantifying the importance of source tasks, thereby enabling an optimal weight assignment that prioritizes source representations with maximal discriminative information and minimal redundancy relative to the target domain.
Furthermore, a recently developed generalization measure based on K-L divergence [59] provides a rigorous foundation for determining optimal transfer quantities. In this work, we demonstrate that such a measure offers a more direct alignment with the cross-entropy loss—a standard objective in machine learning—than previous metrics. Consequently, we extend this metric and framework to jointly optimize source weights and transfer quantities.
@@ III Preliminaries
III-A K-L Divergence Based Measure
We will introduce a K-L divergence based measure as the generalization error measure.
@@ III-A K-L Divergence Based Measure
Compared to other measures, the K-L divergence exhibits a closer correspondence with the generalization error measured by the cross-entropy loss; we formally justify this relationship in Appendix D.
III-B Problem Formulation
A multi-source transfer learning framework consists of a single target task $\mathcal{T}$ and multiple source tasks $\mathcal{S}_{1},\dots,\mathcal{S}_{K}$} that provide auxiliary information for improving performance on the target. To generalize the analysis, we model $\mathcal{T}$ as a parameter estimation problem governed by an underlying distribution $P_{X;\underline{\theta}}$. In typical settings such as supervised classification, this distribution corresponds to the joint probability model over input features $Z$ and labels $Y$, where $X=(Z,Y)$, and $\underline{\theta}$ denotes the parameter of interest. The goal is to accurately estimate the true value of $\underline{\theta}$, which corresponds to optimizing the network parameters of $\mathcal{T}$ in deep learning. Furthermore, we assume that the source tasks and the target task follow the same parametric model and share the same input space $\mathcal{X}$. For notational clarity, we present only the case where $\mathcal{X}$ is discrete, while the theoretical results extend straightforwardly to continuous domains. The target task $\mathcal{T}$} has $N_{0}$ independent and identically distributed (i.i.d.) training samples generated from the underlying distribution $P_{X;{\underline{\theta}}_{0}}$, where the parameter ${\underline{\theta}}_{0}\in\mathbb{R}^{d}$, and we denote the training set as $X^{N_{0}}$. Similarly, the source task $\mathcal{S}_{i}$ has a training set $X^{N_{i}}$ with i.i.d. $N_{i}$ training samples generated from distribution $P_{X;{\underline{\theta}}_{i}}$, where $i\in[1,K]$, and ${\underline{\theta}}_{i}\in\mathbb{R}^{d}$. During training, each sample is weighted in the gradient descent process according to the source weight of its corresponding task. Then, we formulate the training process as a parameter estimation problem and give the form of the estimator. $\hat{{\underline{\theta}}}$ is denoted as the MLE based on the $N_{0}$ samples from $\mathcal{T}$} and $n_{1},\dots,n_{K}$ samples from $\mathcal{S}_{1},\dots,\mathcal{S}_{K}$ with weights $w_{1},\dots,w_{K}$, where $n_{i}\in[0,N_{i}]$ and $w_{i}\in[0,+\infty]$, i.e.,

### v2

@@ Abstract
In multi-source transfer learning, a key challenge lies in how to appropriately differentiate and utilize heterogeneous source tasks. However, existing multi-source methods typically focus on optimizing either the source weights or the amount of transferred samples, largely neglecting their joint consideration. In this work, we propose a theoretical framework, Unified Optimization of Weights and Quantities (UOWQ), that jointly determines the optimal source weights and transfer quantities for each source task. Specifically, the framework formulates multi-source transfer learning as a parameter estimation problem based on an asymptotic analysis of a Kullback–Leibler divergence–based generalization error measure, leading to two main theoretical findings: 1) using all available source samples is always optimal when the weights are properly adjusted; 2) the optimal source weights are characterized by a principled optimization problem whose structure explicitly incorporates the Fisher information, parameter discrepancy, parameter dimensionality, and transfer quantities. Building on the theoretical results, we further propose a practical algorithm for multi-source transfer learning, and extend it to multi-task learning settings where each task simultaneously serves as both a source and a target. Extensive experiments on real-world benchmarks, including DomainNet and Office-Home, demonstrate that UOWQ consistently outperforms strong baselines. The results validate both the theoretical predictions and the practical effectiveness of our framework.
@@ I Introduction
Transfer learning has emerged as a critical paradigm in modern machine learning, especially in scenarios where labeled data is scarce for the target task [56]. By leveraging knowledge from related source tasks, transfer learning enables models to generalize more effectively and efficiently to the target task. In recent years, transfer learning has a wide range of applications. For example, in deep reinforcement learning, transfer learning plays a crucial role in accelerating policy learning and improving sample efficiency by transferring policies, representations, or experience [68]. In healthcare, transfer learning leverages large medical datasets to support rare disease diagnosis with limited clinical data [42]. In natural language processing, pretrained models built on large-scale transfer learning foundations have revolutionized performance across a wide range of downstream tasks [11]. While single-source transfer learning has been studied earlier, growing attention is now being paid to multi-source transfer learning. With the rapid development of large language models (LLMs) trained on massive and diverse corpora, effectively leveraging complementary knowledge from multiple heterogeneous data sources has become increasingly important.
TABLE I: Comparison across matching-based transfer learning methods, based on whether they perform quantity optimization, weight optimization, are tailored to multi-source, have task generality, have shot generality, and require target labels. The ‘✓’ represents obtaining the corresponding aspects, while ‘✗’ represents the opposite. Quantity optimization denotes the ability to determine how many source samples or data subsets should be transferred. Weight optimization denotes the ability to assign or learn weights for different source tasks or samples. Task generality denotes the ability to handle various target task types, and shot generality denotes the ability to avoid negative transfer in different target sample quantity settings, including few-shot and non-few-shot.
@@
In multi-source transfer learning, how to differentially leverage multiple source tasks has emerged as a central research problem. From a statistical perspective, while incorporating additional data from source tasks can reduce estimation variance, sources that are overly dissimilar to the target task may introduce excessive systematic bias, outweighing the benefits of variance reduction. As a result, naive uniform transfer from multiple source tasks can lead to negative transfer [53], making the identification and effective utilization of truly related source tasks a key challenge. Existing approaches typically address this challenge from one of two separate perspectives. On one hand, task- or source-weighting methods aim to control the influence of different sources by assigning importance weights to source models or samples [44, 55], thereby mitigating bias caused by mismatched domains. In existing studies on task weighting, most frameworks implicitly assume that all source samples are utilized, while the transfer weights are determined independently of the transfer quantity. On the other hand, sample selection methods focus on determining the optimal transfer quantity for each source and identifying the source samples to be transferred [18, 23, 60]. Considering source weighting and transfer quantities in isolation restricts the solution space and can lead to suboptimal solutions, whereas jointly optimizing them leads to more principled and effective solutions. The goal of this work is to establish a unified theoretical framework that not only jointly determines the optimal source weights and transfer quantities, but also provides a principled interpretation of their interaction, thereby offering insights into existing weighting and sample selection strategies. In Table I, we provide a detailed comparison of the proposed method with several existing methods.
@@
In this work, we develop a theoretical framework termed Unified Optimization of Weights and Quantities (UOWQ), to determine the optimal weights and quantities for source tasks in transfer learning via asymptotic analysis. Through an asymptotic analysis of the measure based on Kullback–Leibler (K-L) divergence, we formulate the multi-source transfer learning problem as a parameter estimation problem, which in turn yields an optimization problem of the source weights and transfer quantities. First, we prove the conclusion that using all available source samples is always optimal once the weights are properly adjusted, and provide an intuitive interpretation of this result. Moreover, our analysis provides a closed-form solution for the optimal transfer weight in the single-source setting and characterizes the optimal weights in the multi-source setting via a convex optimization formulation, jointly accounting for the Fisher information, parameter discrepancy, parameter dimensionality, and transfer quantities. On this basis, we introduce practical algorithms capable of supporting both multi-source transfer learning and multi-task transfer learning.ă Finally, the effectiveness of the proposed algorithms is validated through experiments on real-world datasets.
@@
  • Theoretical framework. We introduce UOWQ, a mathematical framework based on asymptotic analysis, to jointly optimize transfer quantities and source weights in multi-source transfer learning. Under joint optimization, we prove that using all available source samples is always optimal. In addition, we characterize the optimal transfer weights for each source via an optimization formulation that explicitly accounts for the transfer quantities.
  • Practical algorithm. Based on our theory, we develop practical algorithms derived from the framework for both multi-source transfer learning and multi-task learning. Specifically, to cope with scarce target data and ensure robustness, the algorithms alternate between model optimization and source weight updates.
@@
This work is an extension of our previous conference paper [59], and the additional contributions are summarized as follows. First, the journal version presents a more explicit formulation for jointly optimizing transfer quantities and source weights, together with complete theoretical proofs. In addition, we provide further theoretical analysis, including an intuitive interpretation of why utilizing all source samples remains optimal under joint optimization in Section IV-B, as well as comparisons with related theoretical frameworks in Remark 4 and Table I. Second, in the practical algorithm part in Section V, we further develop a multi-task learning algorithm based on the proposed multi-source transfer learning algorithm. Finally, the experimental section is substantially strengthened with larger-scale datasets, more comprehensive baselines, and evaluations of the proposed multi-task learning extension, along with extensive analyses including ablation studies, weight visualization, robustness evaluation, computational efficiency analysis, and compatibility with Low-Rank Adaptation (LoRA).
The remainder of this paper is organized as follows. Section II reviews the related work. Section III introduces several preliminary results used in our analysis and presents the problem formulation. Section IV develops the main theoretical results to calculate the optimal source weights. Section V introduces practical algorithms. Section VI empirically evaluates the proposed framework and validates the theoretical findings.
@@ II-A Multi-source Transfer Learning
Multi-source transfer learning (MSTL) leverages knowledge from multiple related source domains or tasks to improve performance on a target task, addressing challenges such as domain shift and negative transfer [29, 52, 67]. By emphasizing relevant sources while suppressing less informative ones, MSTL enhances generalization in heterogeneous settings, where effective learning requires balancing source contributions and mitigating distribution discrepancies.
@@ II-B Task weighting
Task weighting, also referred to as source weighting, is a representative class of matching-based approaches in multi-source transfer learning and multi-task learning, as it regulates the contribution of different tasks without explicitly aligning feature distributions. Its objective is to balance competing task objectives during joint optimization. Early approaches relied on static weighting, assigning fixed weights based on heuristics such as task priority or dataset size [2], but these methods lack adaptability to dynamic task interactions. Recent studies have therefore focused on dynamic task weighting strategies. For example, uncertainty-based weighting was introduced in [19], GradNorm balances tasks by normalizing gradient magnitudes [7], and DWA adjusts weights according to task learning speeds [31]. From a multi-objective optimization perspective, MGDA [41] formulates task weighting as a Pareto-optimal optimization problem.
Recent work has further extended task weighting to large-scale pretrained models, primarily through gradient-based matching mechanisms. For instance, task reweighting via gradient alignment has been studied in [66], while conflicting gradients are mitigated by CAGrad [28]. Other studies explore meta-learning–based strategies [45] and task affinity to automate weighting. Despite these advances, task weighting remains challenged by nonstationary task relationships [34], scalability to large task sets, and the lack of strong theoretical guarantees [12, 4]. Benchmark datasets such as Meta-Dataset [48] have been introduced to facilitate standardized evaluation.
Most existing task-weighting methods operate at the model level, assigning scalar weights to source models in a mixture-of-experts fashion [35]. These approaches typically rely on task similarity [32], domain divergence [63], or validation performance [6]. In contrast, only a limited number of studies consider sample-level weighting, where source samples are reweighted and jointly trained with the target data in a unified framework  [44, 62]. Moreover, most existing task-weighting methods implicitly assume access to all source samples, and the resulting weights are independent of the transfer quantity. While recent works have begun to study the problem of optimally selecting transferable source samples or transfer quantities [18, 23, 60], these efforts are largely decoupled from task-weight optimization. Our work provides a unified treatment that jointly optimizes task weights and transfer quantities, enabling principled control over both which sources to transfer from and how much to transfer.
@@ II-C Transfer Learning Theory
Existing theoretical studies in transfer learning can generally be divided into two primary categories. The first category aims to define and quantify the similarity or relatedness between source and target tasks. Numerous metrics have been proposed in this context, including the $l_{2}$ distance [33], optimal transport cost [8], LEEP (Log Expected Empirical Prediction) [37], Wasserstein distance [44], OTCE (Optimal Transport-based Conditional Entropy) [5], LogME [57], NCE[46], GBC (Geometric-Based Correlation) [30], and maximal correlation-based measures [22]. For example, LEEP [37] provides a probabilistic framework for transferability estimation by comparing model predictions on the target task. Meanwhile, OTCE [5] leverages optimal transport theory to quantify transferability under distribution shifts.ă These measures provide principled ways to assess how well knowledge from a source task can potentially benefit a target task.
This work belongs to the second group, which is dedicated to developing theoretical measures that assess and bound the generalization error in transfer learning scenarios. Various generalization error measures have been proposed to guide the assignment of task or source weights, including $f$-divergence [16], mutual information  [1], $\mathcal{H}$-score [55], and ${\mathcal{X}}^{2}$-divergence [47]. Specifically, the $\mathcal{H}$-score analytically characterizes the expected log-loss on the target task under a fixed source feature representation, thereby providing a principled basis for evaluating target predictive performance in transfer learning. Furthermore, a recently developed generalization measure based on K-L divergence [60] provides a rigorous foundation for determining optimal transfer quantities. In this work, we demonstrate that such a measure offers a more direct alignment with the cross-entropy loss—a standard objective in machine learning—than previous metrics. Consequently, we extend this metric and framework to jointly optimize source weights and transfer quantities.
@@ III Preliminaries
In this section, we introduce the theoretical foundations of our framework. We first present the K-L divergence–based generalization error measure in Section III-A, which serves as the optimization objective. We then review the asymptotic normality of the MLE in Section III-B, which enables characterizing estimation variance. Finally, we formulate the multi-source transfer learning problem and its joint optimization objective in Section III-C.
III-A K-L Divergence-Based Measure
We introduce a K-L divergence-based measure as the generalization error measure.
@@ III-A K-L Divergence-Based Measure
Compared to other measures, the K-L divergence exhibits a closer correspondence with the generalization error measured by the cross-entropy loss; we formally justify this relationship in the Appendix.
III-B Asymptotic Normality of the MLE
When attempting to recover the true parameter vector $\underline{\theta}^{*}$ from independent and identically distributed (i.i.d.) observations generated according to the distribution $P_{X;\underline{\theta}^{*}}$, we let $\mathcal{D}$ denote a collection of $n$ such i.i.d. samples. The maximum likelihood estimator (MLE) is then defined as the maximizer of the empirical log-likelihood:ă
Intuitively, the Fisher information matrix characterizes the local geometry of the log-likelihood function with respect to the parameter $\underline{\theta}$. It captures how sensitively the distribution responds to perturbations along different parameter directions, thereby reflecting the statistical identifiability of $\underline{\theta}$.
III-C Problem Formulation
A multi-source transfer learning framework consists of a target task $\mathcal{T}$ and multiple source tasks $\{\mathcal{S}_{1},\dots,\mathcal{S}_{K}\}$ that provide auxiliary information to improve performance on $\mathcal{T}$. To generalize the analysis, we model $\mathcal{T}$ as a parameter estimation problem governed by an underlying distribution $P_{X;\underline{\theta}}$, where $\underline{\theta}$ denotes the parameter of interest and $X$ represents a generic random variable corresponding to the data. For example, in supervised learning, $X=(Z,Y)$ represents the joint distribution over input features $Z$ and labels $Y$. The objective is to accurately estimate the true value of $\underline{\theta}$, which in deep learning corresponds to optimizing the network parameters of $\mathcal{T}$. Furthermore, we assume that the source tasks and the target task follow the same parametric model and share the same input space $\mathcal{X}$. For notational clarity, we present only the case where $\mathcal{X}$ is discrete, while the theoretical results extend straightforwardly to continuous domains.
The target task $\mathcal{T}$ has $N_{0}$ i.i.d. training samples drawn from $P_{X;\underline{\theta}_{0}}$, where $\underline{\theta}_{0}\in\mathbb{R}^{d}$, and the training set is denoted by $X^{N_{0}}$. Similarly, each source task $\mathcal{S}_{i}$ has $N_{i}$ i.i.d. samples from $P_{X;\underline{\theta}_{i}}$, where $i\in[1,K]$ and $\underline{\theta}_{i}\in\mathbb{R}^{d}$, denoted by $X^{N_{i}}$. During training, each sample is weighted in the gradient descent procedure according to the source weight associated with its corresponding task. Motivated by this weighting mechanism, we formulate the training process as a parameter estimation problem and define the corresponding estimator. Specifically, $\hat{{\underline{\theta}}}$ is denoted as the MLE based on the $N_{0}$ samples from $\mathcal{T}$} and $n_{1},\dots,n_{K}$ samples from $\mathcal{S}_{1},\dots,\mathcal{S}_{K}$ with weights $w_{1},\dots,w_{K}$, where $n_{i}\in[0,N_{i}]$ and $w_{i}\in[0,+\infty]$, i.e.,